OpenStudy (anonymous):

if y is inversely proportional; to x, y=6 when x=4 find the constant of proportionality write a formula for y in terms of x and find x when y=8

OpenStudy (anonymous):

y is inversely proportional to x therefore xy=k= constant ....for all values of x and y. For y=6 then x=4 so k=(6)(4)=24 the equation becomes xy=24 i.e y=24/x At y=8 8=24/x so, x=24/8 x=3

OpenStudy (anonymous):

okay thanks alot..lifesaver i need this for my exam 2morrow